Circular No. 8966

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 Q1 (MATIČIČ) A moving object of stellar appearance discovered by Stanislav Matičič

on four 40-s R-band CCD images taken with the 60-cm f/3.3 Cichocki robotic telescope in strong moonlight in the course of the Comet and Asteroid Search Program at Črni Vrh Observatory (discovery observation tabulated below) has been found by other observers to show cometary appearance. H. Mikuž adds that a series of 1-min unfiltered exposures obtained with the same telescope around Aug. 19.816 UT under darker sky conditions show that the object has a somewhat-diffuse appearance (coma diameter ∼ 10′′) with strong condensation. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD observers have also noted the object’s cometary appearance. R. Ligustri (Talmassons, Udine, Italy, 0.35-m f/5 reflector, Aug. 19.9) reports that twenty 120-s co-added images show a short tail elongation ∼ 18′′ long. R. Holmes and H. Devore (Charleston, IL, U.S.A., 0.81-m f/4.0 astrograph, Aug. 20.07–20.08) report that the object appears diffuse with a wide tail extending 37′′.3 in p.a. 87^{o}.9. K. Sárneczky (0.60-m Schmidt telescope at Konkoly Observatory, Aug. 20.8, 360-s coadded image) writes that the comet shows a 8′′ coma and a tail 10′′ long in p.a. 120^{o}. J. M. Aymami (Tiana, Spain, 0.25-m f/4.3 Schmidt-Cassegrain reflector, Aug. 20.8) notes that co-added images show a distinctly nebulous object with slight oval shape (∼ 9′′ long along its major axis). V. Gonano, E. Guido, and G. Sostero (Remanzacco, Italy, 0.45-m f/4.4 reflector, Aug. 20.8) write that sixteen 60-s co-added unfiltered exposures show a coma diameter of ∼ 10′′ and a broad tail 15′′ long toward p.a. 122^{o}. S. Foglia (Novara Veveri, Italy, 0.40-m f/4 reflector, Aug. 20.8) reports that images obtained by P. Concari, D. Crespi, G. Galli, O. Lesca, S. Minuto, and himself show the comet to be diffuse with a 10′′ coma.

2008 UT  α_{2000}  δ_{2000}  Mag.
Aug. 18.81364  19^{h}02^{m}50.^{s}20  +50^{o}51′19′′.5  17.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-Q12.

T = 2009 Jan. 2.546 TT  ω = 105.^{o}645
Ω = 9.460 }2000.0
q = 2.94139 AU  i = 118.672
2008 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green